# Understanding the Issue Types in Open Source Blockchain-based Software Projects with the Transformer-based BERTopic

### v1

@@ Abstract
Blockchain-based software systems are increasingly deployed across diverse domains, yet a systematic understanding of their development challenges remains limited. This paper presents a large-scale empirical study of 497,742 issues mined from 1,209 open-source blockchain projects hosted on GitHub. Employing BERTopic, a transformer-based topic modeling technique, we identify 49 distinct issue topics and organize them hierarchically into 11 major subcategories. Our analysis reveals that both general software development issues and blockchain-specific concerns are nearly equally represented, with Wallet Management and UI Enhancement emerging as the most prominent topics. We further examine the temporal evolution of issue categories and resolution times, finding that Wallet issues not only dominate in frequency but also exhibit the longest resolution time. Conversely, Mechanisms issues are resolved significantly faster. Issue frequency surged after 2016 with the rise of Ethereum and decentralized applications, but declined after 2022. These findings enhance our understanding of blockchain software maintenance, informing the development of specialized tools and practices to improve robustness and maintainability.
Index Terms:
Blockchain, Github Issues, Topic Modeling, BERTopic, Resolution Time
@@ I Introduction
Recent research has focused on identifying issues and bugs in blockchain-based systems, aiming to uncover the underlying characteristics and trends that developers encounter. However, these studies have either focused narrowly on bugs within a limited number of blockchain projects [11] or examined other types of development activities, such as commits and pull requests [12]. To the best of our knowledge, no prior research has undertaken a systematic, large-scale analysis of issue reports from blockchain-based applications on GitHub to comprehensively characterize the development challenges prevalent in this domain.
@@ II Background & Related Works
Several recent studies have applied topic modeling with GitHub issue analysis to study blockchain-related software development. Vaccargiu et al. [22] utilized BERTopic on GitHub issues and comments to uncover sustainability-related themes in Ethereum developer discussions, focusing on the Go-Ethereum repository. In a related study, Vaccargiu and Tonelli [23] applied BERTopic to blockchain projects in the environmental sector, identifying core concerns such as carbon credit tracking and resource optimization. Wan et al. [11] manually classified 946 unique bugs across eight blockchain systems into ten categories using card sorting. They also investigated bug-fixing dynamics in blockchain systems, revealing that security bugs took the longest, while performance bugs often remained unresolved for extended periods.
Despite extensive work on issue analysis, topic modeling, and blockchain software, there remains a gap in applying advanced topic modeling techniques to systematically uncover and categorize development challenges in blockchain-related GitHub issues. Existing work typically targets general software domains or specific concerns, lacking a comprehensive, data-driven view of developer-reported issues in blockchain projects. This study fills that gap.
@@ III Methodology
The methodology is organized into two principal phases: (i) data collection and preprocessing of GitHub repositories and issues, and (ii) topic modeling using BERTopic to uncover recurring topics. Each phase is described in detail below.
@@ III-A1 Project Selection
First, for repositories created before January 2021, we used the curated dataset from Das et al.[12], based on a validated GHTorrent subset [25]. This dataset originally included 1,194 repositories filtered by cryptocurrency relevance, platform affiliation, and popularity. However, as their dataset represents a snapshot up to January 2021, many of these repositories are found to be inactive or archived. Applying additional activity-based filtering, we retained 218 active repositories for analysis.
@@
Additional automated filtering was applied to reduce noise: repositories were excluded if they were empty, archived, inactive in the past year, or had fewer than five stars or forks. These criteria, based on empirical guidelines [12, 24], aimed to balance dataset inclusiveness with quality. Inactivity was used as a proxy for abandonment, while stars and forks indicated public interest and collaboration. After filtering, 6,833 repositories remained.
@@ III-B Topic Modeling with BERTopic
Figure 2: BERTopic [20] can be viewed as a sequence of 5 steps to create its topic representations. The first three steps are for creating the topics, and the last three steps are for representing and interpreting the topics.
@@
BERTopic begins by converting documents into embeddings using transformer models. Unlike traditional topic modeling, which requires extensive data preprocessing [30, 19], it requires minimal preprocessing. As noted in the official BERTopic documentation [^7] https://maartengr.github.io/BERTopic/faq.html?#should-i-preprocess-the-data, traditional preprocessing is unnecessary when using document embeddings, as every part of a document contributes to understanding its overall topic. But, it is recommended to clean the text of excessive noise, such as HTML tags, as these do not contain contextual information. Following these, we only removed markup tags, emojis, and URLs from the data.
@@
We initially used concatenated issue titles and bodies for topic modeling, but this led to noisy, semantically diffuse clusters with thousands of topics. Analysis revealed that many bodies were auto-generated by GitHub Actions or CI/CD tools or template-based, offering little useful content—an issue also noted in prior studies [32, 33]. To improve topic quality and interpretability, we limited modeling to issue titles, which are typically concise and semantically rich.
After generating an embedding, the next step is to cluster the embeddings. To facilitate faster clustering, we applied Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction. This choice is justified by UMAP’s ability to preserve global semantic relationships while enhancing local cluster separability [34]. We retained 100 dimensions and set the neighborhood size to 25, balancing computational efficiency with contextual richness in high-volume data.
@@ IV Approach, Analysis & Results
After generating the topics, the next critical step involved assigning semantically meaningful and representative labels to each topic to facilitate interpretation. This labeling was performed collaboratively by the first and second authors, both researchers with expertise in software engineering and testing, through an iterative process of discussion and refinement. For each topic, 30 highly representative issues were examined—ranked based on their probability values—to determine the most suitable label. The process continued until both researchers reached a consensus. Figure 3 illustrates the distribution of issue percentages across all 49 identified topics. The top five most prominent topics identified across all issues were: Wallet Management and Connectivity (21.3%), UI Enhancement (10.6%), Logging and Error handling (9.2%), CICD (4.6%), and RPC and Networking (4.5%). The high volume of UI-related issues likely reflects the complexity of blockchain technology and unintuitive interfaces. Even popular wallets pose challenges: new users struggle with complex interactions [37], while experienced users face unfamiliar metaphors and transaction flows [38]. Our analysis supports this, with frequent terms like icon, modal, and mobile in issues.
After completing the labeling, the topics were further grouped into higher-level subcategories and categories by merging semantically and conceptually related topics into a hierarchical structure, guided by domain knowledge and following practices observed in prior research [39, 40]. For example, topics such as Write, Fix, Update Tests and E2E Test were placed under the broader subcategory of Testing. This categorization resulted in two overarching topics: Software Development Issues (53.87%) comprising 4 primary subcategories, and Blockchain-Specific Issues (46.13%) consisting of 7 primary subcategories. The hierarchical structure of these topics is illustrated in Figure 4.
@@
To examine how resolution times vary across subcategories, we computed the minimum, maximum, mean, and median durations for each subcategory (Table I). These statistics reveal notable differences in resolution patterns across issue types. For example, some issues are closed within minutes of creation. Instances like MystenLabs/sui/issues/1180888https://github.com/MystenLabs/sui/issues/1180 and Vagabonds-Labs/cofiblocks/issues/100 demonstrate this behavior. Prior studies have reported such cases where developers write a patch, commit it to the version control system (VCS), and subsequently create and immediately close the corresponding issue [11, 8, 15].
@@
The Wallet subcategory exhibits the longest maximum resolution time of 7.9 years (2,896 days), followed by categories like Infrastructure and Transaction, which also have long extremes. On the other hand, Mechanisms shows the shortest maximum resolution time at 2.8 years (1,022 days), though most subcategories range between 4 and 7 years. These findings are consistent with previous work, which attributes long resolution times to poor issue descriptions, unclear reproducibility, or shifting project priorities [11, 8, 15].
The mean resolution time across subcategories generally ranges from approximately 20 to 46 days. Mechanisms issues have the shortest mean resolution time (20 days), indicating quicker turnarounds, while the Wallet category again stands out with the longest average (46 days), suggesting more complex or resource-intensive problem domains. Other categories, such as Security & Data Integrity and Transaction, also report higher means (37–45 days).
Regarding median resolution time, values range from just 0.6 to 2.9 days, further highlighting how many issues are resolved swiftly. The shortest median is seen in Mechanisms (0.62 days), and the longest in Security & Data Integrity (2.93 days). This skew, with a small median but relatively large mean, suggests the presence of outliers—most issues are resolved quickly, but a few take significantly longer.
TABLE I: Resolution Time (in days) of Closed Issues by Subcategory, Sorted by Mean Resolution Time.
Category Min Max Mean Median
Mechanisms 0.000035 1022.12 19.81 0.62
Development & Maintenance 0.000012 2201.83 20.50 0.70
Documentation 0.0 1595.29 20.69 0.60
User Experience & Interaction 0.000012 2537.03 27.50 0.87
Infrastructure 0.0 2218.66 28.82 0.98
Community Interaction 0.000046 1935.54 32.00 1.21
Networking 0.000035 2319.35 32.34 0.96
Digital Asset 0.0 2057.94 32.55 1.14
Transaction 0.000035 2474.69 36.65 1.80
Security & Data Integrity 0.000035 2446.68 45.47 2.93
Wallet 0.000023 2896.19 45.95 2.02
@@
To understand how issue categories evolved over time, we analyzed quarterly issue counts for each of the 11 main subcategories from 2012 to 2024. Figure 6 visualizes these trends, capturing the trajectory of reported issues across the blockchain ecosystem.
@@
A significant rise in issue frequency is seen across nearly all subcategories beginning in 2016. A similar trend has also been observed in previous studies on dApps [41, 42]. This sharp increase coincides with the launch and growing adoption of platforms like Ethereum, which officially released its Homestead version in March 2016. Ethereum introduced a new paradigm of decentralized applications (dApps) and smart contracts, significantly expanding the scope of blockchain software development. This growth spurred the creation of a vast number of blockchain projects, which in turn generated an increasing number of reported issues, particularly in categories like Development & Maintenance, User Experience & Interaction, and Wallet.
From 2023 onward, we observed a noticeable decline in issue frequency across most subcategories. This trend may stem from ecosystem maturation, better documentation, and reusable tools that reduce the need for extensive issue tracking, though further research is needed. Despite the decline, subcategories like Wallet and UI continue to exhibit high activity, reflecting their central role in user interaction and asset management.
@@
The frequency of issues surged after 2016, coinciding with the rise of Ethereum and dApps, and then declined after 2022. Wallet, the most frequently reported subcategory, consistently remained the most reported throughout the timeline.
@@ V Discussion
This study provides a large-scale empirical analysis of blockchain software projects using GitHub issue mining and BERTopic modeling. Analysis identifies the most common issue topics in blockchain projects and shows that Development & Maintenance and Wallet-related issues are most prominent. The dominance and prolonged resolution of wallet issues reflect their inherent complexity. This signals a pressing need for specialized tools, standard frameworks, and secure design practices tailored to wallet functionalities.
Despite longer resolution times for Wallet and Security issues, around 80% of issues are resolved within a month, indicating that most are relatively manageable. This suggests practitioners should allocate resources based on issue complexity, while researchers have an opportunity to develop blockchain-specific tools to address more demanding issues.
Trend analysis shows that the issue frequency rose after 2016 with Ethereum and dApps, then declined after 2022. Wallet issues remained prominent, with persistent UI and Infrastructure concerns. These trends show how platform evolution shapes development, urging practitioners to adapt and researchers to pursue longitudinal studies on evolving practices and tools.
Overall, this work reinforces the notion that blockchain software engineering extends beyond conventional development practices, requiring tailored methodologies, domain-specific tooling, and a deeper understanding of decentralized environments. For the research community, the findings offer a foundation for future studies aimed at mitigating the unique challenges of blockchain systems. For practitioners—such as software company leadership—the findings suggest focusing resources on high-issue areas like Wallet Management and Connectivity. By recognizing these domain-specific needs, both communities can contribute to more secure, maintainable, and user-friendly blockchain applications.
V-A Limitations

### v2

@@ Abstract
Blockchain-based software systems are increasingly deployed across diverse domains, yet a systematic understanding of their development challenges remains limited. This paper presents a large-scale empirical study of 497,742 issues mined from 1,209 open-source blockchain projects hosted on GitHub. Employing BERTopic, a transformer-based topic modeling technique, we identify 49 distinct issue topics and organize them hierarchically into 11 major subcategories. Our analysis reveals that both general software development issues and blockchain-specific concerns are nearly equally represented, with Wallet Management and UI Enhancement emerging as the most prominent topics. We further examine the temporal evolution of issue categories and resolution times, finding that Wallet issues not only dominate in frequency but also exhibit the longest resolution time. Conversely, Mechanisms issues are resolved significantly faster. Issue frequency surged after 2016 with the rise of Ethereum and decentralized applications, but started declining after 2022. These findings enhance our understanding of blockchain software maintenance, informing the development of specialized tools and practices to improve robustness and maintainability.
@@ I Introduction
Recent research has focused on identifying issues and bugs in blockchain-based systems, aiming to uncover the underlying characteristics and trends that developers encounter. However, these studies have either focused narrowly on bugs within a limited number of projects [11] or examined other types of development activities, such as commits and pull requests [12]. To the best of our knowledge, no prior research has undertaken a systematic, large-scale analysis of issue reports from blockchain-based applications on GitHub to comprehensively characterize the development challenges in this domain.
@@ II Background & Related Works
Several recent studies have applied topic modeling with GitHub issue analysis to study blockchain-related software development. Vaccargiu et al. [22] utilized BERTopic on GitHub issues and comments to uncover sustainability-related themes in Ethereum developer discussions. In a related study, Vaccargiu and Tonelli [23] applied BERTopic to blockchain projects in the environmental sector, identifying core concerns such as carbon credit tracking and resource optimization. Wan et al. [11] manually classified 946 unique bugs across eight blockchain systems into ten categories using card sorting and investigated bug-fixing dynamics in blockchain systems. Furthermore, Wan et al. [24] analyzed blockchain-related discussions from Stack Overflow. Using LDA, they extracted 45 topics distributed across six architectural layers. Although this research shares objectives with ours, it diverges in its reliance on Q&A platforms as the data source. Repository mining on GitHub, by contrast, offers a more direct lens into the processes of software development and evolution, as issue reports typically capture long-term maintenance challenges [25], whereas Stack Overflow posts predominantly reflect short-term problem solving and knowledge sharing [26]. This difference between these two data sources motivated us to choose GitHub in our study.
Despite substantial progress in issue analysis, topic modeling, and blockchain software research, advanced topic modeling has not yet been systematically applied to GitHub issues in the blockchain domain. Prior studies either address general software contexts or focus on specific concerns, leaving a gap in providing a comprehensive, data-driven view of developer-reported challenges. This study seeks to address that gap.
@@ III Methodology
The methodology is organized into two principal phases: (i) data collection and preprocessing of GitHub repositories and issues, (ii) topic modeling using BERTopic to uncover recurring topics. Each phase is described in detail below.
@@ III-A1 Project Selection
First, for repositories created before January 2021, we used the curated dataset from Das et al.[12], based on a validated GHTorrent subset [27]. This dataset originally included 1,194 repositories filtered by cryptocurrency relevance, platform affiliation, and popularity. However, as their dataset represents a snapshot up to January 2021, many of these repositories are found to be inactive or archived. Applying additional filtering to remove such projects, we retained 218 active repositories.
@@
Additional automated filtering was applied to reduce noise: repositories were excluded if empty, archived, inactive in the past year, or had fewer than five stars or forks. These criteria, based on empirical guidelines [12, 25], aimed to balance dataset inclusiveness and quality. Inactivity served as a proxy for abandonment, while stars and forks indicated interest and collaboration. After filtering, 6,833 repositories remained.
@@ III-B Topic Modeling with BERTopic
Figure 2: BERTopic [20] can be viewed as a sequence of 6 steps to create its topic representations. The first three steps are for creating the topics, and the last three steps are for representing and interpreting the topics.
@@
BERTopic begins by converting texts into embeddings using transformer models. Unlike traditional topic modeling, which requires extensive data preprocessing [32, 19], it requires minimal preprocessing. As noted in the official BERTopic documentation [^7] https://maartengr.github.io/BERTopic/faq.html?#should-i-preprocess-the-data, traditional preprocessing is unnecessary when using embeddings, as every part of a document contributes to understanding its overall topic. But, it is recommended to clean the text of excessive noise, such as HTML tags, as these do not contain contextual information. Following this, we removed only markup tags, emojis, and URLs from the data.
@@
We initially used concatenated issue titles and bodies for topic modeling, but this led to noisy, semantically diffused clusters with thousands of topics. Analysis revealed that many bodies were auto-generated by GitHub Actions or CI/CD tools or template-based, offering little useful content—an issue also noted in prior studies [34, 35]. To improve topic quality and interpretability, we limited modeling to issue titles, which are typically concise and semantically rich.
After generating an embedding, the next step is to cluster the embeddings. To facilitate faster clustering, we applied Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction. This choice is justified by UMAP’s ability to preserve global semantic relationships while enhancing local cluster separability [36]. We retained 100 dimensions and set the neighborhood size to 25, balancing computational efficiency with contextual richness in high-volume data. These parameters were chosen through trial and error, as the combination of high data volume and numerous parameters made a grid-search type approach challenging.
@@ IV Approach, Analysis & Results
After generating the topics, the next critical step involved assigning semantically meaningful and representative labels to each topic to facilitate interpretation. This labeling was performed collaboratively in real-time by the first and second authors, both researchers with 3 and 8 years of expertise in software engineering and testing, through an iterative process of discussion and refinement. For each topic, 30 highly representative issues were examined, ranked based on their probability values, to determine the most suitable label. The process continued until both researchers reached a consensus. Figure 3 illustrates the distribution of issue percentages across all 49 identified topics. The top five most prominent topics identified across all issues were: Wallet Management and Connectivity (21.3%), UI Enhancement (10.6%), Logging and Error handling (9.2%), CICD (4.6%), and RPC and Networking (4.5%). The high volume of UI-related issues likely reflects the complexity of blockchain technology and unintuitive interfaces. Even popular wallets pose challenges: new users struggle with complex interactions [39], while experienced users face unfamiliar metaphors and transaction flows [40]. Our analysis supports this, with frequent terms like icon, modal, and mobile in issues.
After labeling, the topics were recursively aggregated into higher-level subcategories by merging semantically and conceptually related topics, yielding 11 subcategories at the intermediate level and two categories at the root of the hierarchical structure. This process was guided by domain knowledge and aligned with practices reported in prior research [41, 42]. For example, topics such as Write, Fix, Update Tests and E2E Test were placed under the broader subcategory of Testing. This categorization resulted in two overarching categories: Software Development Issues (53.87%) comprising 4 primary subcategories, and Blockchain-Specific Issues (46.13%) consisting of 7 primary subcategories. The hierarchical structure of these topics is illustrated in Figure 4.
@@
To examine how resolution times vary across subcategories, we computed the minimum, maximum, mean, median, and $10^{th}$ and $90^{th}$ percentile durations for each subcategory (Table I). These statistics reveal notable differences in resolution patterns across issue types. For example, some issues are closed within minutes of creation. Instances like MystenLabs/sui/issues/1180888https://github.com/MystenLabs/sui/issues/1180 and Vagabonds-Labs/cofiblocks/issues/100 demonstrate this behavior. Prior studies have reported such cases where developers write a patch, commit it to the version control system, and then create and immediately close the corresponding issue [11, 8, 15].
@@
TABLE I: Resolution time (in days) of closed issues by subcategory. Reported values include minimum, maximum, mean, median, and the $10^{\text{th}}$ and $90^{\text{th}}$ percentiles. Sorted by mean time.
Category Min Max Mean Median $P_{10}$ $P_{90}$
Mechanisms 0.000035 1022.12 19.81 0.62 0.0231 17.74
Development & Maintenance 0.000012 2201.83 20.50 0.70 0.00279 35.01
Documentation 0.0 1595.29 20.69 0.60 0.00161 29.89
User Experience & Interaction 0.000012 2537.03 27.50 0.87 0.00375 46.84
Infrastructure 0.0 2218.66 28.82 0.98 0.00632 52.36
Community Interaction 0.000046 1935.54 32.00 1.21 0.01012 71.79
Networking 0.000035 2319.35 32.34 0.96 0.00589 67.80
Digital Asset 0.0 2057.94 32.55 1.14 0.00786 59.80
Transaction 0.000035 2474.69 36.65 1.80 0.01126 76.33
Security & Data Integrity 0.000035 2446.68 45.47 2.93 0.01337 117.87
Wallet 0.000023 2896.19 45.95 2.02 0.00856 98.06
Resolution times vary widely across subcategories, from minutes to several years. The longest maximum resolution time occurs in the Wallet category (7.9 years), while the shortest is in Mechanisms (2.8 years), with most subcategories ranging from 4 to 7 years. Mean resolution times span from 19.8 days (Mechanisms) to 46 days (Wallet), indicating that issues in categories such as Wallet, Security & Data Integrity, and Transaction are more resource-intensive or complex. Median resolution times, however, remain lower, ranging from 0.6 days (Documentation) to 2.9 days (Security & Data Integrity), showing that 50% of issues are addressed promptly, while a small share of long-standing cases skew the mean. Such prolonged cases are consistent with prior findings linking delays to vague issue descriptions, limited reproducibility, or shifting project priorities [11, 8, 15]. The percentile data further illustrates this skew. At the 10th percentile, resolution times are consistently below one hour across all subcategories. By contrast, the 90th percentile ranges from 17.7 days (Mechanisms) to 118 days (Security & Data Integrity).
@@
To understand how issue categories evolved, we analyzed quarterly issue counts for each of the 11 subcategories from 2012 to 2024. Figure 6 visualizes these trends, capturing the trajectory of reported issues across the blockchain ecosystem.
@@
A significant rise in issue frequency is seen across nearly all subcategories beginning in 2016. A similar trend has also been observed in previous studies on dApps [43, 44]. This sharp increase coincides with the launch and adoption of platforms like Ethereum, which officially released its Homestead version in March 2016. Ethereum introduced a new paradigm of decentralized applications (dApps) and smart contracts, significantly expanding the scope of blockchain software development. This growth spurred the creation of many blockchain projects, which in turn generated an increasing number of reported issues, particularly in categories like Development & Maintenance, User Experience & Interaction, and Wallet.
From 2023 onward, we observed a noticeable decline in issue frequency across most subcategories. This trend may stem from ecosystem maturation, better documentation, reusable tools that reduce the need for extensive issue tracking, or a broader shift of open-source efforts toward AI-related projects, though further research is needed. Despite the decline, subcategories like Wallet and UI continue to exhibit high activity, reflecting their role in user interaction and asset management.
@@
The frequency of issues surged after 2016, coinciding with the rise of Ethereum and dApps, then started declining after 2022. Wallet, the most frequently reported subcategory, consistently remained the most reported throughout the timeline.
@@ V Discussion
This study provides a large-scale empirical analysis of open-source blockchain software projects using GitHub issue mining and BERTopic modeling. Findings show that Development & Maintenance and Wallet-related issues dominate, with nearly half being blockchain-specific. Resolution time analysis reveals that Security and Wallet issues take the longest to resolve, making them particularly resource-demanding. Trend analysis further indicates that issue frequency grew during the rise of Ethereum and dApps after 2016 but started declining after 2022.
The findings have important implications for both research and practice. First, the predominance of blockchain-specific issues (nearly 50%) demonstrates that conventional software engineering practices alone cannot adequately address the unique challenges of decentralized systems. Practitioners must adopt blockchain-tailored approaches to issue tracking and triaging. In particular, the longer resolution times in max and 90th percentile for Security and Wallet issues highlight the need for dedicated resources, stronger expertise, and specialized frameworks. By prioritizing based on complexity, project maintainers can improve efficiency and reduce bottlenecks.
Second, the dominance of Wallet issues underscores their technical and security complexity. Wallets lie at the intersection of usability, interoperability, and security, making them vulnerable to long-standing challenges. This suggests a pressing need for standardized frameworks, automated testing strategies, and reusable tools to streamline wallet development.
Third, the temporal analysis illustrates how platform innovations shape issue spaces. The surge of issues after 2016 and the decline after 2022 suggest that issue patterns reflect ecosystem shifts, likely influenced by the introduction of Ethereum and decentralized applications, and later by a global focus shift toward artificial intelligence. Researchers should therefore pursue longitudinal studies to examine how practices, tools, and issue categories evolve in response to such shifts.
Finally, the emergence of novel issue types, such as governance mechanisms and community interaction, indicates that blockchain development will continue diversifying rather than stabilizing. This dynamic nature makes continuous developer training essential, ensuring practitioners are prepared for new application domains and problem spaces.
Overall, this study shows that blockchain projects need methodologies, tooling, and expertise beyond traditional practices. For practitioners, the findings suggest prioritizing resources on Wallet Management, Security, and Infrastructure. For researchers, they offer a foundation for developing specialized tools and longitudinal studies. Addressing these needs can improve the security, maintainability, and usability of blockchain systems as they expand into new domains.
V-A Limitations & Future Work